### v2

# From Rankings to Insights: Evaluation Should Shift Focus from Leaderboard to Feedback

Abstract
Automatic evaluation benchmarks such as MT-Bench, Arena-Hard, and Auto-Arena are seeing growing adoption for the evaluation of Large Language Models (LLMs). Existing research has primarily focused on approximating human-based model rankings using limited data and LLM-as-a-Judge. However, the fundamental premise of these studies, which attempts to replicate human rankings, is flawed. Specifically, these benchmarks typically offer only overall scores, limiting their utility to leaderboard rankings, rather than providing feedback that can guide model optimization and support model profiling. Therefore, we advocate for an evaluation paradigm shift from approximating human-based model rankings to providing feedback with analytical value. To this end, we introduce Feedbacker, an evaluation framework that provides comprehensive and fine-grained results, thereby enabling thorough identification of a model’s specific strengths and weaknesses. Such feedback not only supports the targeted optimization of the model but also enhances the understanding of its behavior. Feedbacker comprises three key components: an extensible tree-based query taxonomy builder, an automated query synthesis scheme, and a suite of visualization and analysis tools. Furthermore, we propose a novel LLM-as-a-Judge method: PC2 (Pre-Comparison-derived Criteria) pointwise evaluation. This method derives evaluation criteria by pre-comparing the differences between several auxiliary responses, achieving the accuracy of pairwise evaluation while maintaining the time complexity of pointwise evaluation. Finally, leveraging the evaluation results of 17 mainstream LLMs, we demonstrate the usage of Feedbacker and highlight its effectiveness and potential. Our project homepage and dataset are available at https://liudan193.github.io/Feedbacker/.
1 Introduction
Evaluating large language models (LLMs) has become increasingly important as these models advance in various capabilities [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Among them, Chatbot Arena [14], which uses over 2.8 million human-labeled cases, provides relatively accurate evaluations for 223 LLMs (as of 2025-04-09). However, its labor-intensive nature limits its applicability, which is particularly pronounced during model development, where timely feedback is essential. These limitations have motivated researchers to explore automatic evaluation [15, 16, 17, 18, 19, 8], which uses a small amount of data and LLM-as-a-Judge [9, 20, 21, 22, 23] to approximate the model rankings in Chatbot Arena. Automatic evaluation benchmarks not only accelerate leaderboard updates but also enable timely feedback for LLM development.
Challenge 1. However, although automatic evaluations can provide timely feedback, they are often insufficient to guide model optimization or profiling. This is primarily due to the fundamental premise of approximating final Chatbot Arena rankings, which introduces several critical limitations. First, to fit given rankings with limited data, these benchmarks typically employ a restricted query set that excessively prioritizes mainstream tasks. This narrow scope neglects long-tail queries reflective of real-world use cases [15, 16, 17, 18, 19, 8], thereby overlooking some critical aspects of model capabilities. Second, this premise lacks the requirement for fine-grained taxonomic annotations, which restricts these benchmarks to generating only overall scores for rankings. This limitation prevents identifying specific weaknesses and strengths, limiting the benchmarks’ analytical value.
Figure 1: An overview of Feedbacker framework.
Solution 1. To tackle these limitations and make automatic evaluation more suitable for generating feedback, we introduce Feedbacker, an extensible evaluation framework built on TaxBuilder and RealMix. TaxBuilder constructs a query-type taxonomy, which is crucial for providing a solid foundation for comprehensive and fine-grained evaluation. The taxonomy includes first-level nodes representing domains (e.g., writing, coding), second-level nodes for classification principles (e.g., writing style, task type), and third-level nodes and beyond for fine-grained tags. TaxBuilder is extensible, offering features including node insertion, refinement, pruning, and layer pruning to support future expansion. RealMix generates high-quality and realistic queries from a set of seed data, simulating real users, preventing data contamination, and further enhancing extensibility.
Challenge 2. Compared to existing benchmarks, our evaluation dataset presents a unique challenge. Since we aim for a comprehensive evaluation, the overall dataset size is relatively large, yet the number of samples per specific query type remains small to allow for fine-grained analysis. This characteristic poses a significant challenge for existing LLM-as-a-Judge paradigms. Evaluation paradigms typically fall into three categories: pairwise evaluation [19], pointwise evaluation [24, 25], and win rate based on pairwise evaluation [16, 2, 26, 18]. In our evaluation context, pairwise evaluation is impractical due to its quadratic growth in time cost as the number of models increases, despite offering superior accuracy [15]. Pointwise evaluation is more efficient but sacrifices accuracy [15, 4, 27]. Additionally, the win rate metric demands a relatively large dataset to ensure reliable results [16]. Therefore, we pose a critical question: Can an LLM-as-a-Judge method achieve the accuracy of pairwise evaluation, the efficiency of pointwise evaluation, and reliability with limited data?
Solution 2. To address this question, we propose PC2 pointwise evaluation. It begins by generating query-specific criteria, which are then used to assess responses. Its core innovation lies in leveraging responses generated by a diverse set of models and using the differences among these responses to establish evaluation criteria. Since the criteria for each query can be fixed, they can be reused when evaluating new models, achieving an evaluation time similar to pointwise evaluation.
In summary, our main contributions are four-fold:
$\bullet$ We propose shifting evaluation paradigm from rankings to feedback. To this end, we introduce Feedbacker, which employs comprehensive and fine-grained queries to identify specific strengths and weaknesses. The feedback generated by Feedbacker can facilitate model optimization and profiling.
$\bullet$ Utilizing Feedbacker, we develop the initial version of the taxonomy and evaluation dataset for six domains: writing, roleplay, knowledge, coding, mathematics, and reasoning. These are denoted as Feedbacker-T-V0 (Fig. 2) and Feedbacker-D-V0 (Tab. 4), respectively.
$\bullet$ We introduce PC2 pointwise evaluation, which achieves higher accuracy than pairwise evaluation, maintains efficiency of pointwise evaluation, and ensures reliability with limited evaluation dataset.
$\bullet$ To showcase the usage of Feedbacker, we evaluate 17 mainstream LLMs and conduct in-depth analyses of the results. By uncovering several fine-grained insights, we demonstrate the effectiveness and potential of Feedbacker.
2 Feedbacker
The overview of Feedbacker is shown in Fig. 1. It consists of three components: TaxBuilder for building an extensible query taxonomy, RealMix for generating new queries based on the taxonomy, and visualization and analysis toolkits for user-friendly interaction.
2.1 TaxBuilder: Automatic Tree-based Taxonomy Building
Figure 2: A subset of Feedbacker-T-V0. Please refer to project page for complete taxonomy.
To establish a comprehensive and fine-grained evaluation system, we first need to develop the query taxonomy. To achieve this, we introduce TaxBuilder, an automatic and extensible solution for constructing tree-based taxonomy from large volumes of unstructured queries. The workflow of TaxBuilder is illustrated in Fig. 3.
Preparation. TaxBuilder initiates the process with a manually constructed basic taxonomy tree $T_{init}$. This basic tree is simple and requires minimal human effort (refer to § C.1.1 for details). Then, we employ a low-cost model (gpt-4o [28]) to annotate query tags for a large set of real user queries  [^1] We utilize the dataset made available by the Chatbot Arena team, accessible at https://www.kaggle.com/competitions/wsdm-cup-multilingual-chatbot-arena/data.. The primary aim of this phase is to generate a large pool of unstructured query tags, named $TG_{init}$. Further details regarding this step are elaborated in § C.1.2.
Figure 3: An overview of TaxBuilder, a tree-based automatic taxonomy generation.
Node Insertion. The most core design of TaxBuilder is its node insertion mechanism. A naive approach would be to input the entire taxonomy as context to a powerful LLM and ask it to identify the position for the new node. However, this task involves both long-context and complex reasoning, which is challenging even for the most advanced LLMs. Moreover, as the taxonomy grows, the input context becomes increasingly longer and the reasoning more difficult, thereby limiting the extensibility of this approach.
To address this issue, we introduce a recursive node insertion process inspired by how trees are constructed in data structures. When a new node needs to be inserted, we traverse the tree step by step. At each step, the LLM only needs to determine the relationship between the new node and the nodes at the current level. In this way, the model only considers a small portion of the tree and reasons over three options: (1) the node already exists, (2) it is a sibling of the current nodes, or (3) it should be added as a child of one of them. This design keeps the context short and the reasoning simple, making TaxBuilder significantly more scalable as the taxonomy expands.
Formally, given a node $tg$ and the current tree $T_{cur}$ (we represent the tree by its root), we invoke LLM-as-Decision-Maker, denoted as $DM_{ins}$, as follows:
$DM_{ins}(tg,T^{c}_{cur})\rightarrow\{\text{<E>},\text{<S>},T^{c_{i}}_{cur}\},$ (1)
where $T^{c}_{cur}=\{T^{c_{1}}_{cur},T^{c_{2}}_{cur},\ldots,T^{c_{m}}_{cur}\}$ denotes the set of first-level child nodes of $T_{cur}$. The function of $DM_{ins}$ is to determine the relationship between $tg$ and $T^{c}_{cur}$. The prompt for $DM_{ins}$ is presented in Fig. 15. It will return one of three possible relationship types:
  • <E> (Exists): This indicates that the node $tg$ already exists in $T^{c}_{cur}$ or has a semantically equivalent node. In this case, the new node $n$ is discarded to avoid redundancy.
  • <S> (Sibling): This signifies that $tg$ is a sibling of the nodes $T^{c}_{cur}$. Consequently, $T^{c}_{cur}$ is updated to include $tg$, i.e., $T^{c}_{cur}=\{T^{c_{1}}_{cur},T^{c_{2}}_{cur},\ldots,T^{c_{m}}_{cur}\}\cup\{tg\}$.
  • $\mathbf{T^{c_{i}}_{cur}}$ (Child Node): This implies that $tg$ should be a child of the specific node $T^{c_{i}}_{cur}$. In this scenario, a recursive insertion is performed.
After recursively traversing the entire tree, we successfully insert the node $tg$ and obtain the tree for the next iteration. We iteratively perform this operation for all nodes, and ultimately obtain a tree that contains all tags from $TG_{init}$.
Node Refinement and Pruning. Node refinement serves two purposes: (1) resolving incorrect parent-child relationships between nodes at the same level, and (2) splitting leaf nodes that conflate multiple concepts into separate, more atomic concepts (e.g., separating "science fiction fanfiction" into "science fiction" and "fanfiction"). Node pruning addresses two issues: (1) merging duplicate nodes at the same level, and (2) removing rarely encountered concepts that unnecessarily complicate the tree. We perform the refinement and pruning processes recursively, guided by the LLM-as-Decision-Maker (see Fig. 16), which effectively corrects faulty nodes and controls node complexity.
Layer Pruning. We introduce two rules to reduce tree depth: (1) Meaningless depth pruning: if a parent node has only a single child, the child is pruned to eliminate unnecessary depth; (2) Complexity pruning: the tree is restricted to a maximum of four layers. Nodes beyond the fourth layer are retained only if they have five or more child nodes, as such structures are usually meaningful. which effectively controls tree complexity and prevents overgrowth.
After completing node insertion, node refinement, node pruning, and layer pruning, we obtain the final taxonomy, which we call Feedbacker-T-V0. The current taxonomy tree is easy to interpret by human, allowing for convenient manual modification if needed. A simplified version of this tree is shown in Fig. 2. The formal algorithm for the tree construction process is provided in Alg. 1.
2.2 RealMix: New Query Synthesis
We opt for synthetic data over existing datasets for two reasons: (1) existing datasets may not contain enough queries with the required tags in taxonomy, and (2) to prevent data contamination. To ensure the evaluation data includes high-quality, realistic queries that can be encountered in real-world, we propose RealMix. RealMix operates by mixing several appropriate content details from multiple real-user queries and adapting them to align with the specified tags. While there are other data synthesis approaches [29, 30] that mix multiple queries, none are capable of generating queries with specified tags. The core advantage of RealMix in fact lies in that it is built upon the TaxBuilder, which enables controlled and tag-aligned synthesis.
Seed Data. We first use QwQ-32b [31, 32] to annotate a set of real-user queries [33] with three types of labels: domain, query tags, and quality. The domain label covers six domains: writing, roleplay, knowledge, coding, mathematics, and reasoning. Query tags are assigned based on Feedbacker-T-V0. For quality annotation, each domain is paired with a checklist of seven criteria; queries meeting at least four are marked as high-quality. In total, we collect approximately 31,000 high-quality annotated queries. Further details are provided in § C.2.1.
Query Synthesis. RealMix leverages content from real user queries to generate queries aligned with specific tags. This approach allows us to obtain realistic queries associated with desired tags. To synthesize a query, we sample one reference query and three content queries. A query generation model then extracts appropriate real-world content from the content queries and generates a new query that incorporates these contents while matching the tags of the reference query. If no suitable content is found for the specified tags, this generation is discarded. We adopt reference queries instead of random tag combinations from Feedbacker-T-V0 since some combinations do not exist in practice. To mitigate model bias, we employ multiple query generation models (deepseek-v3 [34], gpt-4o, and doubao-1.5pro [35]). More details are provided in § C.2.2.
Verification and Filtration. We validate both the tags and the quality of each generated query using the same procedure applied to the seed data. Then, we discard the low-quality queries. To further reduce model bias, for each query, we randomly sample checking models for quality and tags from three reasoning models (QwQ-32B, deepseek-r1 [36], and o1-mini [37]).
2.3 Visualization and Analysis Toolkits
We provide a suite of visualization and analysis tools to assist users in understanding and interpreting evaluation results. Due to space constraints, the description of each tool is provided in § C.3.
3 Dataset Description
Dataset Statistics. We construct Feedbacker-D-V0, the initial version of the evaluation dataset, which consists of 3,343 queries, each annotated with a domain and tags. Tab. 4 summarizes the key statistics of Feedbacker-D-V0. We ensure that the sample size for each tag in Feedbacker-T-V0 is at least 19 (#MinNum) to guarantee the statistical representativeness of the evaluation results.
Table 1: Statistics of Feedbacker-D-V0.
Domain #Samples #Tags #MinNum Avg.Length
Writing 1108 594 19 772.57
Roleplay 470 429 19 1008.46
Knowledge 540 315 20 608.57
Coding 636 369 19 1232.03
Mathematics 344 189 20 817.02
Reasoning 245 186 19 904.81
Total 3343 2082 19 880.92
Figure 4: Comparison of RealMix and real user queries, judged by five human evaluators.
Quality Control. We validate data quality through three tasks involving human and automatic evaluation: Five graduate-level annotators perform two comparative assessments. For each generated query and its corresponding real user query (i.e., reference query), annotators evaluate (1) which query is higher in quality and (2) which is more likely to occur in real-world scenarios. Results in Fig. 4 show generated queries are consistently preferred, confirming the effectiveness of RealMix. Details are in § C.4.1. (3) We assess potential data leakage in generated queries in § C.4.2.
4 Pre-Comparison-derived Criteria Pointwise Evaluation
In this section, we first formulate traditional evaluation paradigms and highlight their strengths and limitations in Sec. 4.1. Then, we present the motivation and description of our method in Sec. 4.2.
4.1 Problem Formulation
Pairwise Evaluation. Pairwise evaluation [38, 39] refers to comparing two responses to determine which is better. Given two LLM responses $y^{A}$ and $y^{B}$ for a given instruction $x$, a judge model $J$ is tasked to provide a comparative assessment with a pairwise evaluation prompt $p^{I}$.
$J(y^{A},y^{B}|x,p_{i})\rightarrow\{A>B,A=B,A<B\}.$ (2)
Pairwise evaluation allows for capturing relative performance, which is often more reliable than pointwise evaluation. However, as the number of models evaluated increases, the time consumption grows significantly [15, 4], with a quadratic complexity of O(n2).
Vanilla Pointwise Evaluation. Vanilla pointwise evaluation [40, 41] refers to scoring each response independently. The judge model $J$ is tasked to assign a score to a single response with a vanilla pointwise evaluation prompt $p_{o}$. Typically, a baseline response $y_{b}$ is used to help stabilize evaluation standards.
$J(y|x,p_{o},y_{b})\rightarrow s\in\mathbb{R}.$ (3)
Pointwise evaluation is straightforward and does not incur unacceptable evaluation costs as the number of models being evaluated increases. However, it often produces inferior evaluation results compared to pairwise evaluation [4, 15].
4.2 PC2-Pointwise Evaluation
Motivation. Our core insight is that, given the superior performance of pairwise evaluation over pointwise evaluation, we can adapt the mechanism of pairwise evaluation into the pointwise setting to create an enhanced pointwise evaluation. Pairwise evaluation excels in comparing two responses by identifying their differences, evaluating which response performs better with respect to each difference, and ultimately selecting the winner based on advantages in the most important differences. (See examples in § D.) In contrast, pointwise evaluation struggles to identify such differences because it evaluates each response independently. To bridge this gap, we introduce a pre-comparison phase, where we leverage several auxiliary responses to extract relevant evaluation criteria before scoring. These criteria act as proxies for the key differences identified in pairwise evaluation. At the same time, we prompt the model to assign weights to each criterion, emulating how pairwise evaluation prioritizes more significant differences.
Pre-Comparison-derived Criteria Extraction. The core design of PC2 pointwise evaluation lies in extracting criteria from a pre-comparison process. Specifically, given an instruction $x$, we prompt $n$ LLMs $\{M_{1},M_{2},...,M_{n}\}$ to produce $n$ responses $\{y^{1},y^{2},...,y^{n}\}$. The selection of multiple LLMs is intended to diversify the responses, thereby facilitating the extraction of comprehensive, diverse and effective criteria. Subsequently, we assign the judge model $J$ two tasks. The first task involves using $J$ to compare the $n$ responses, thereby surfacing their differences (i.e., criteria). This is done using prompt $p^{c}$. The second task is to weight these criteria using prompt $p_{w}$:
$J(\{y^{1},y^{2},...,y^{n}\}\mid x,p_{c},p_{w})\rightarrow C=\{c_{1},c_{2},...,c_{m}\},W=\{w_{1},w_{2},...,w_{m}\}$
where $C$ represents the criteria, and $W$ denotes the corresponding weights, where $\sum_{i=1}^{m}w_{i}=100$. Since we use a reasoning LLM as the judge model, we typically assign two tasks to the model simultaneously. If the judge model is a chat model, we recommend handling these tasks separately.
Evaluation Execution. Once the criteria are determined, we can score a single response $y$:
$J(y|x,p_{c},y_{b},C,W)\rightarrow s\in\mathbb{R},$ (4)
where $p_{c}$ represents our criteria-weighted evaluation prompt and $s$ is the final score.
This approach leverages the comprehensive understanding gained through pre-comparison to establish query-specific criteria, enabling more effective evaluation while maintaining computational efficiency as the number of evaluated models increases.
5 Results on PC2 Pointwise Evaluation
In this section, we conduct the experiments on different LLM-as-a-Judge methods. The experiment details are shown in § B.1.
5.1 Main Results
Table 2: Comparison of accuracy across different LLM-as-a-Judge methods.
Method Overall Writing Roleplay Knowledge Mathematics Coding Reasoning
Pairwise Evaluation 0.8176 0.8000 0.8302 0.8716 0.8750 0.8160 0.7672
Vanilla Pointwise Evaluation 0.6116 0.5838 0.7358 0.5229 0.7708 0.6960 0.5259
   + Baseline Answer 0.6525 0.6054 0.6038 0.6881 0.7500 0.6800 0.6466
NCD Pointwise Evaluation (No Pre-Com) 0.6688 0.6378 0.7358 0.6389 0.7917 0.7097 0.6207
PC2 Pointwise Evaluation (Single Model Pre-Com) 0.6792 0.6432 0.7358 0.6697 0.7500 0.6880 0.6810
   + Diverse Model Pre-Com 0.7453 0.7351 0.8491 0.7248 0.7500 0.7440 0.7328
    + Baseline Answer 0.8318 0.7892 0.7925 0.8257 0.8542 0.8720 0.8707
Main Results. As demonstrated in Tab. 2, we can observe the following: (1) Base pairwise evaluation yields superior performance than vanilla pointwise evaluation, which is consistent with previous study [15]; (2) Vanilla pointwise evaluation performs poorly, but introducing a baseline answer significantly improves accuracy. (3) The naive criteria decomposition (NCD) pointwise evaluation outperforms pointwise evaluation, indicating that performing criterion decomposition alone can significantly enhance performance. (4) Introducing a single model for pre-comparison yields only a marginal improvement in performance compared to NCD, indicating that pre-comparison significantly enhances the quality of the extracted metrics. (5) When diverse models are adopted for pre-comparison, there is a significant improvement, especially in open-domain tasks such as writing and roleplay. (6) Further incorporating baseline answer leads to additional significant improvement, even surpassing the pairwise evaluation. Results on more benchmarks are shown in § A.1.
(a) Point scale.
(b) Criteria number.
Figure 5: Ablation study.
Figure 6: Consistency of rankings across 17 models with varying sample sizes.
Ablation Study. As shown in Fig. 6, we observe slight fluctuations in performance across different point scales. Notably, the model demonstrates superior performance when the point scale is smaller. This is intuitive, as multiple evaluations tend to yield more stable scores with a small point scale. As for number of criteria, too few (e.g., 2-3) or too many (e.g., 3-15) criteria can both adversely affect performance. Based on these findings, we use a 1-5 point scale and 3-9 criteria in all experiments.
Determining the Minimum Meaningful Sample Size. If an evaluation dataset has only one sample for the HTML category, can it reliably indicate which models perform well or poorly in this category? Obviously, different samples can lead to inconsistent evaluation results. This raises an important question: What is the minimum number of samples needed per query type for consistent evaluation? To address this, we analyze tags with over 32 samples. For each tag, we randomly select $N$ samples 20 times, rank 17 LLMs based on each sample set, and measure consistency using the average Spearman’s rho coefficient across the 20 rankings. We then average the consistency scores across all tags. As shown in Fig. 6, consistency improves with larger sample sizes. Our Feedbacker-D-V0 dataset contains at least 19 samples per tag, consistently achieving a score above 0.9, which can be considered reliable.
Figure 7: Token cost comparison.
Table 3: Agreement rate on the winner between two similarly performing models with a small data size.
Model Pair Avg. Min. Std. Dev.
qwen-max-2024-10-15 vs. doubao-1-5-pro-32k-250115
ours 81.52% 45.00% 0.1264
win rate 52.88% 15.00% 0.1586
Meta-Llama-3.1-70B-Instruct vs. Qwen2.5-72B-Instruct
ours 89.39% 55.00% 0.1260
win rate 70.76% 5.00% 0.1875
Meta-Llama-3.1-8B-Instruct vs. Qwen2.5-7B-Instruct
ours 81.67% 35.00% 0.1695
win rate 49.24% 20.00% 0.1801
5.2 Comparison with Other Evaluation Methods
Comparison with Pairwise Evaluation. We compare output token cost with pairwise evaluation and our method. From the results in Fig. 7, we observe that as the number of models increases, the token cost burden of pairwise evaluation becomes significant, whereas our method exhibits linear growth. This indicates that our method is more scalable for comparing a larger number of LLMs.
Comparison with Win Rate. To assess the effectiveness of the win rate metric under limited data conditions, we select pairs of models with similar performance and focus on query types over 100 instances. For each query type, we use the evaluation results from all available data as the ground-truth winner. We then perform 20 random samplings, each consisting of 19 instances, and compare the winner obtained from the sampled instances with the ground-truth winner. The average agreement rate is calculated across all query types and samplings. As shown in Tab. 3, the win rate metric performs poorly with small datasets, highlighting its limitations when data is scarce and model differences are subtle. In contrast, our evaluation method provides a more consistent evaluation.
6 Results on Feedbacker
6.1 A Quick Overview of the Evaluation Results.
To utilize Feedbacker for analysis, we first employ Feedbacker-D-V0 to evaluate 17 mainstream LLMs in this section. The experimental setup is provided in § B.3. An overview of evaluation results is presented in § B.4. For complete results, please refer to the project homepage.
6.2 Exploring Feedbacker: Practical Usage Examples
(a) Rankings of 17 LLMs under Markup Languages.
(b) Fine-grained rankings.
Figure 8: Model rankings for root.coding.Programming Languages.Markup Languages.
In this section, we showcase the usage of Feedbacker to analyze an LLM’s strengths and weaknesses by several examples. These examples demonstrate the effectiveness and potential of Feedbacker.
To begin with, the Automatic Weakness Reporter reveals an interesting anomaly: gpt-4o-2024-11-20 ranked 1st under the fine-grained query tag root.coding.Programming Languages.Markup Languages.XML, while its overall ranking in Tab. 5 is 6th. This suggests an unusual strength in that specific category.
To further explore this phenomenon, we examine the XML’s parent node root.coding.Programming Languages.Markup Languages using the Visualization Tool in the “By Query” view, which displays model rankings across its subdomains. As shown in Fig. 8(a), claude3.7-sonnet-20250219 ranks 1st overall, while gpt-4o-2024-11-20 ranks 5th. This result indicates that gpt-4o-2024-11-20 does not perform as strongly across the entire Markup Languages category—its lead in XML may be an outlier or the result of targeted optimization.
We then adopt the Failure Mode Explorer to get the results in Fig. 8(b), which reveals a notable distinction: claude3.7-sonnet-20250219 maintains consistent high performance across all subdomains (low standard deviation), whereas gpt-4o-2024-11-20’s strong ranking is mainly driven by its exceptional performance in XML. Its scores in other subdomains are relatively low. This indicates that gpt-4o-2024-11-20 may have been specially optimized for XML tasks, but does not exhibit a broad advantage across the entire Markup Languages domain. In contrast, claude3.7-sonnet-20250219 demonstrates stable, and well-rounded performance. This not only exposes a weakness in gpt-4o-2024-11-20, but also suggests a need for broader optimization.
7 Related Work
Automatic Evaluation Benchmark. We review existing benchmarks here, focusing on their motivations, dataset designs, and evaluation methods. AlpacaEval-2.0 [16] uses 805 curated general-domain data points to assess instruction-following with win rate metric. MT-Bench [15] includes 80 high-quality queries across 8 domains. It adopts pointwise evaluation on a 1–10 scale. Arena-Hard [18] filters 500 challenging Chatbot Arena datasets and uses win rate metric for evaluation. LiveBench [3] manually updates queries in real-time with rule-based scoring to reduce LLM biases. Auto-Arena [19] employs pairwise debates for Elo-based ranking, with a 40-query evaluation set. WildBench [2] updates real user queries to reflect real-world distributions, using both win rate and pointwise evaluation. We discuss the differences between our methods and existing studies in § A.1.
8 Conclusion
We propose shifting the focus of evaluation from only rankings and leaderboard to providing valueable feedback. To this end, we introduce Feedbacker, an extensible evaluation framework featuring comprehensive and fine-grained queries, along with tools for visualization and analysis. Our proposed PC2 pointwise evaluation achieves superior accuracy compared to pairwise evaluation while retaining pointwise efficiency. This approach effectively balances precision and speed, offering a practical solution for LLM evaluation. We demonstrate that utilizing Feedbacker enables the convenient identification of a model’s strengths and weaknesses, thereby providing valuable feedback.
While our work employs comprehensive, fine-grained queries to deliver feedback, this approach is not the sole option. We believe Feedbacker can serve as a starting point to inspire the community to develop diverse, impactful feedback mechanisms, enhancing the analytical value of evaluations.
We conduct more in-depth discussions including (1) the implications for future research in § A.2, and (2) the limitations of our work in § A.3.
Appendix A More Discussion
A.1 Distinctions from Prior Works Enabled by Fine-Grained Dataset
Several existing studies, such as SELF-INSTRUCT [42], Nemotron-4-340B [43], WildBench [2], CLEVA [44], and C2LEVA [5], provide fine-grained evaluation of LLM capabilities. However, their taxonomies are relatively coarse and lack extensibility. For instance, C2LEVA classifies LLM abilities into four broad categories—Language, Knowledge, Reasoning, and Harms—and further subdivides Language into only two types: Typo-Fixing and Transliteration. In contrast, our benchmark introduces an unprecedentedly detailed, fine-grained, and extensible taxonomy. Additionally, none of the aforementioned benchmarks provide visualization or analytical tools tailored for fine-grained evaluation, which is a key contribution of our work.
A concurrent study, EvalTree [45], also organizes queries into a tree structure. However, our TaxBuilder differs fundamentally. (1) Unlike EvalTree, which reorganizes existing datasets, our method starts with a fixed, extensible taxonomy and automatically synthesizes data to populate it. (2) TaxBuilder supports multi-dimensional classification based on topic, style, domain, and programming language, whereas EvalTree organizes queries solely by model capability. (3) Our method fully leverages the power of advanced reasoning models, resulting in more accurate classifications. In contrast, EvalTree relies on sentence embeddings—which may lack precision—and K-Means clustering, with cluster labels generated by LLMs. (4) We provide analysis tools to identify model strengths and weaknesses, which are absent in EvalTree. (6) We also addresses sample size requirements for reliable evaluation within subcategories, an important factor not considered by EvalTree.
A.2 Implications for Future Research
In the future, we believe Feedbacker will provide valuable insights to the community.
(1) For model training, Feedbacker can help identify a model’s strengths and weaknesses. By quantifying fine-grained performance, it enables developers to adjust training parameters accordingly, such as data distribution.
(2) For model evaluation, Feedbacker prevents overestimating model capabilities. Its broad coverage across diverse categories allows it to reveal that models excelling in mainstream tasks may still struggle significantly in less common or specialized areas.
(3) For data synthesis, to the best of our knowledge, TaxBuilder offers an unprecedentedly scalable taxonomy that empowers developers to generate comprehensive, diverse and complex synthetic data for model training. For synthetic data, a concurrent work focusing on constructing code feature trees for code generation tasks [46] conveys ideas similar to those of TaxBuilder.
A.3 Limitations of Our Work
Currently, our framework is limited to language models and does not yet support multimodal models [47, 48, 49, 50, 51]. In future work, we plan to introduce a multimodal extension, Feedbacker-Anything. Moreover, our framework currently lacks support for evaluations related to AI mechanisms such as safety, honesty, factuality, hallucination, and fairness [52, 53, 54, 55, 56, 57, 57, 58, 59, 60, 61]. However, our approach is highly extensible and can be easily adapted to include them as new evaluation domains.
Appendix B Experiments
B.1 Experimental Setup for Evaluation Method
Models. The P2C pointwise evaluation involves two types of model. In the pre-comparison stage, multiple models are utilized to generate responses. Here we employ gpt-4o-2024-11-20 [28], doubao-1-5-pro-32k-250115 [35], and deepseek-v3-250324 [34] to ensure the diversity of the responses. For judge model, we adopt QwQ-32B [31] due to its lightweight nature and superior reasoning performance. Unless otherwise specified, we use the officially recommended decoding parameters.
Dataset. All our experiments are based on pairwise human preference datasets Feedbacker-HPD and RMB [62]. Pairwise human preference datasets usually contain a prompt and two responses, and the model’s task is to determine which response is better. For pairwise evaluation, the judge model directly selects the better response. For pointwise evaluation, the judge model assigns scores to both responses, with the higher-scoring response being chosen as the winner.
Feedbacker-HPD is a high-quality human preference dataset containing 636 samples. Queries are carefully filtered, and responses are annotated with majority human labels, making the dataset more representative of human preferences. A detailed description of Feedbacker-HPD is provided in § C.4.4. RMB is a recently popular large-scale human preference dataset comprising 10K queries for helpfulness and 7K for harmlessness. We only use the helpfulness subset in paper.
Unless otherwise specified, our experiments are conducted using our own constructed dataset, Feedbacker-HPD, which will be publicly released alongside Feedbacker-D-V0. Results on RMB are presented only in § B.2.
Baseline. We employ pairwise evaluation, vanilla pointwise evaluation, and naive criteria decomposition (NCD) as baselines. NCD refers to generating query-specific criteria based only on the query, without the process of pre-comparison [63, 7, 64]. Additionally, we also tested the impact of various components on P2C pointwise evaluation. "Single Model Pre-Com" refers to using only gpt-4o-2024-11-20 with a temperature of 1, generating three responses. "Diverse Model Pre-Com" involves using gpt-4o-2024-11-20, doubao-1-5-pro-32k-250115, and deepseek-v3-250324 as models for generating auxiliary responses. For the baseline answer, we use gpt-4o-2024-11-20 with a temperature of 1.0 and top-p set to 0.95, generating a single response as the baseline answer. Regardless of whether we use vanilla pointwise evaluation or our method, when adding the baseline answer, we first evaluate it using the method without the baseline answer and include the evaluation result in the prompt. This process helps standardize the evaluation. Detailed instructions can be found in the respective prompts.
Prompt. For PC2 pointwise evaluation, the prompt for pre-comparison-derived criteria extraction is shown in Fig. 9, the prompt for evaluation execution is shown in Fig. 10, the prompt for evaluation execution with baseline answer is shown in Fig. 11. For pairwise evaluation and vanilla pointwise evaluation, we directly use the prompts from MT-Bench [15]. For NCD, the corresponding prompt is displayed in Fig. 12.
PC2 Pointwise Evaluation (Pre-Comparison-derived Criteria Extraction) You are an impartial judge responsible for evaluating the quality of responses provided by different LLMs to a given [question]. Your task is to design a comprehensive evaluation framework that includes clearly defined metrics and their respective weights. You shuold answer step by step. You should answer in English. Please carefully follow these steps:
1. **Analyze Responses**: You must first compare several provided [answers] and identify their differences. The objective of this comparison is to pinpoint distinguishing factors that significantly influence the quality of the responses.
2. **Develop Metrics**: Establish a hierarchical set of evaluation metrics. There should be 3 to 9 primary metrics. Each primary metric should have several detailed sub-metrics to provide specific, measurable criteria for evaluating the responses.
3. **Assign Weights**: Allocate appropriate weights to each metric based on its relative importance in distinguishing the quality of the responses. The weights should be integers, and the sum of all weights should equal 100.
4. **Output Format**: Present the final evaluation framework in a structured list format. You do not need to include the primary metrics; only the secondary metrics are required, in the following format:
<Evaluation_Framework>
1. Description of Secondary Metric 1 | Weight 1
2. Description of Secondary Metric 2 | Weight 2
3. Description of Secondary Metric 3 | Weight 3
<Evaluation_Framework>
[User Question]
{question}
[The Start of Assistant 1’s Answer]
{answer_1}
[The End of Assistant 1’s Answer]
[The Start of Assistant 2’s Answer]
{answer_2}
[The End of Assistant 2’s Answer]
[The Start of Assistant 3’s Answer]
{answer_3}
[The End of Assistant 3’s Answer]
Figure 9: Prompt for pre-comparison-derived criteria extraction.
PC2 Pointwise Evaluation (Evaluation Execution) Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider the evaluation system displayed below. Begin your evaluation by providing an explanation for each metric, assessing the response objectively. Score each metric on a scale of 1 to 3, where 1 represents complete failure to meet the criterion and 3 represents perfection. The weights of the metrics must sum to 100, and the final weighted score should be calculated on a scale up to 300, reflecting the weighted sum of the individual scores. After evaluation, you must summarize the results within <The Start of Evaluation Result> and <The End of Evaluation Result>. Below is an example output:
<The Start of Evaluation Result>
Metric 1 | score: [2]
Metric 2 | score: [3]
Metric 3 | score: [1]
Final Weighted Score: [[200]]
<The End of Evaluation Result>
[Question]
{question}
[The Start of Evaluation System]
{eval_system}
[The End of Evaluation System]
[The Start of Assistant’s Answer]
{answer}
[The End of Assistant’s Answer]
Figure 10: Prompt for evaluation execution.
PC2 Pointwise Evaluation (Evaluation Execution with Baseline Answer) Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. Your evaluation should consider the evaluation system displayed below. Begin your evaluation by providing an explanation for each metric, assessing the response objectively. Score each metric on a scale of 1 to 3. Use the baseline answer as a baseline; a higher score indicates a better response compared to the baseline answer, while a lower score indicates a worse response. The weights of the metrics must sum to 100, and the final weighted score should be calculated on a scale of 100 to 300, reflecting the weighted sum of the individual scores. After evaluation, you must summarize the results within <The Start of Evaluation Result> and <The End of Evaluation Result>. Below is an example output:
<The Start of Evaluation Result>
Metric 1 | score: [2]
Metric 2 | score: [3]
Metric 3 | score: [1]
Final Weighted Score: [[200]]
<The End of Evaluation Result>
[Question]
{question}
[The Start of Evaluation System]
{eval_system}
[The End of Evaluation System]
[The Start of Baseline Answer]
{answer_baseline}
[The End of Baseline Answer]
[The Start of Evaluation for Baseline Answer]
{critic_baseline}
[The End of Evaluation for Baseline Answer]
[The Start of Assistant’s Answer]
{answer}
[The End of Assistant’s Answer]
Figure 11: Prompt for evaluation execution with baseline answer.
Prompt of Naive Criteria Decomposition Pointwise Evaluation You are an impartial judge responsible for evaluating the quality of responses provided by different LLMs to a given [question]. Your task is to design a comprehensive evaluation framework that includes clearly defined metrics and their respective weights. You shuold answer step by step. You should answer in English. Please carefully follow these steps:
1. **Analyze Question**: You must first analyze the question. The objective of this comparison is to find important factors that significantly influence the quality of the responses.
2. **Develop Metrics**: Establish a hierarchical set of evaluation metrics. There should be 3 to 9 primary metrics. Each primary metric should have several detailed sub-metrics to provide specific, measurable criteria for evaluating the responses.
3. **Assign Weights**: Allocate appropriate weights to each metric based on its relative importance in distinguishing the quality of the responses. The weights should be integers, and the sum of all weights should equal 100.
4. **Output Format**: Present the final evaluation framework in a structured list format. You do not need to include the primary metrics; only the secondary metrics are required, in the following format:
<Evaluation_Framework>
1. Description of Secondary Metric 1 | Weight 1
2. Description of Secondary Metric 2 | Weight 2
3. Description of Secondary Metric 3 | Weight 3
<Evaluation_Framework>
[User Question]
{question}
Figure 12: Prompt for naive criteria decomposition.
B.2 More Experiment on PC2-Pointwise Evaluation
In this section, we present the accuracy of PC2 pointwise evaluation on the RMB [62] benchmark. The results shown in Tab. 4 demonstrate that while PC2 maintains the pointwise evaluation paradigm, it achieves results comparable to pairwise evaluation.
Table 4: Comparison of accuracy across different LLM-as-a-Judge methods on RMB [62].
Method Overall Generation Chat Role Playing Brainstorming Summarization Rewrite Classification Translation Open QA Closed QA Code Reasoning
Pairwise Evaluation 0.7427 0.8000 0.6566 0.7949 0.7469 0.7544 0.7374 0.8333 0.7267 0.7014 0.6923 0.8280 0.7331
Vanilla Pointwise Evaluation 0.6197 0.6723 0.5060 0.7436 0.5768 0.6316 0.5354 0.7222 0.6333 0.6082 0.5692 0.7640 0.5657
PC2 Pointwise Evaluation 0.7353 0.8000 0.6687 0.8333 0.7676 0.6667 0.6667 0.7778 0.6867 0.7315 0.6718 0.7600 0.7689
   + Baseline Answer 0.7394 0.7940 0.7108 0.7821 0.7083 0.6842 0.6263 0.6667 0.6846 0.7890 0.7077 0.7800 0.7490
B.3 Experimental Setup for Evaluation and Leaderboard
Models. We select 17 mainstream LLMs for evaluation. For closed-source models, we select claude3.7-sonnet-20250219 [65], gpt-4o-2024-11-20 [28], qwen-max-2024-10-15 , doubao-1-5-pro-32k-250115 [35], and hunyuan-standard-2025-02-10 [66]. For open-source models, we select deepseek-v3-250324 [34], deepseek-r1-250120 [36], QwQ-32B [31], Gemma-3-27B-IT [67], Qwen2.5-72B-Instruct [68], Gemma-3-4B-IT [67], Qwen2.5-32B-Instruct [68], Meta-Llama-3.1-70B-Instruct [69], Qwen2.5-7B-Instruct [68], Meta-Llama-3.1-8B-Instruct [69], Mistral-7B-Instruct-v0.3 , and Phi-4-Mini-Instruct [70]. Among these, deepseek-v3-250324 and deepseek-r1-250120 are deployed with fp8, while the other models are deployed with bf16. For all models, if their official decoding parameters are provided, we use the default settings. If not, we set the temperature to 0.7 and top_p to 0.95.
B.4 Leaderboard
We propose in the paper how to quickly check the weaknesses of a LLM through the tree structure. Due to the limitation of the paper’s length, we only present the first-level nodes (i.e., the six domain) in the paper, while the more fine-grained results are presented on our project website. The leaderboard results, as presented in Tab. 5, reveal some valuable weaknesses of existing LLMs. These weaknesses can significantly guide the model’s development if they are timely available during the development process of the model. Deepseek-v3-250324 ranks at the forefront among all tested models due to its superior coding, mathematics, and reasoning capabilities. However, it is noteworthy that its roleplay ability exhibits significant deficiencies. Claude3.7-sonnet-20250219 demonstrates high proficiency in coding, yet its overall capabilities are relatively weak. Gemma-3-4B-IT achieves exceptionally high scores in liberal arts subjects, but it performs poorly in coding and mathematics, which are more science-oriented. Additionally, it is worth mentioning that doubao-1-5-pro-32k-250115 possesses a notable advantage in mathematical abilities.
Table 5: Overall and domain-level scores (fine-grained results on project page).
Model Size Overall Writing Roleplay Knowledge Coding Mathematics Reasoning
Closed-source and Open-source (>100B) LLMs
deepseek-v3-250324 671B 454.47(1) 454.08(3) 443.75(5) 464.54(2) 452.89(1) 450.28(1) 464.53(1)
deepseek-r1-250120 671B 453.76(2) 468.82(1) 461.19(1) 460.32(4) 434.31(3) 418.86(3) 456.35(4)
[…]
Open-source (<100B) LLMs
QwQ-32B 32B 452.75(3) 465.50(2) 455.78(2) 464.94(1) 426.15(4) 430.91(2) 462.08(3)
Gemma-3-27B-IT 27B 437.61(4) 449.74(4) 452.91(3) 462.22(3) 399.08(6) 392.74(4) 462.13(2)
Qwen2.5-72B-Instruct 72B 395.04(7) 406.01(8) 378.90(9) 409.28(8) 383.07(7) 372.91(8) 407.08(9)
Gemma-3-4B-IT 4B 391.92(9) 420.14(7) 431.55(6) 426.82(7) 326.42(12) 292.43(14) 421.03(7)
Qwen2.5-32B-Instruct 32B 370.52(11) 386.67(12) 356.79(10) 380.65(11) 342.44(10) 360.14(10) 388.93(11)
Meta-Llama-3.1-70B-Instruct 70B 362.23(12) 388.10(11) 360.94(9) 368.04(12) 336.79(11) 305.13(13) 381.05(12)
Qwen2.5-7B-Instruct 7B 357.59(13) 378.48(13) 348.52(12) 367.58(13) 324.79(13) 339.03(11) 369.78(13)
Meta-Llama-3.1-8B-Instruct 8B 338.07(15) 366.84(14) 345.72(13) 345.94(15) 310.33(15) 271.33(16) 341.63(16)
Mistral-7B-Instruct-v0.3 7B 314.22(16) 351.71(16) 331.22(15) 324.97(16) 272.39(16) 223.40(17) 324.49(17)
Phi-4-Mini-Instruct 3.8B 309.55(17) 328.02(17) 320.75(16) 307.78(17) 268.48(17) 287.93(15) 345.47(15)
Appendix C Taxonomy and Dataset
C.1 TaxBuilder
C.1.1 Manually Initialized Basic Taxonomy Tree
TaxBuilder requires a manually crafted initial taxonomy as a starting point. This step is quite straightforward. First, we need to define a root node—typically a domain—such as coding.
Next, to capture the various aspects of this domain, TaxBuilder asks users to manually define several classification principles, such as task types, technical domains, and programming languages.
Finally, we manually construct a basic fine-grained taxonomy. Fig. 13 shows an example of a manually created taxonomy for the coding domain. Additional examples can be found in our open-source code repository.
Figure 13: Initial taxonomy for coding domain.
C.1.2 Initial Annotation for Query Type Tags
We use gpt-4o-2024-11-20 to generate a large number of low-quality query type tags. The prompt used for generation is shown in Fig. 14.
Initial low-quality Type Labeling ### Task
Your task is to categorize a given query based on the following rules. For each rule, explain your reasoning before assigning a label. Evaluate the query against each domain in the order provided below until a match is found. Once a domain is identified, proceed with its specific sub-rules. If no domain matches, proceed to the next domain check. If no specific domain applies, return a generic "other" label.
#### Rules
##### Step 1: Determine the Domain
Evaluate the query against the following domains in this order. Stop and proceed with the corresponding sub-rules once a domain is matched. If the domain is labeled "other", return {"domain": "other", "domain_name": "<custom_domain_name>"} and do not proceed further. Explain your reasoning before assigning a label.
1. Writing - Condition: If the query relates to writing, purposeful professional writing, literature, storytelling, grammar, or language-related topics. - Label: "writing" - If not matched, proceed to the next domain.
2. Roleplay - Condition: If the query involves roleplay, character interactions, storytelling, or immersive scenario-based dialogue. - Label: "roleplay" - If not matched, proceed to the next domain. 3. Coding - Condition: If the query relates to programming, algorithms, or code-related topics. - Label: "coding" - If not matched, proceed to the next domain. 4. Mathematics - Condition: If the query pertains to mathematical concepts, computations, proofs, or problem-solving. - Label: "mathematics" - If not matched, proceed to the next domain. 5. Reasoning - Condition: If the query pertains to reasoning, logic, critical thinking, or problem-solving without direct reference to specific knowledge or programming. - Label: "reasoning" - If not matched, proceed to the next domain. 6. Knowledge - Condition: If the query pertains to factual knowledge or general subject areas such as science, history, literature, philosophy, or current affairs. - Label: "knowledge" - If not matched, label as "other" with a custom "domain_name" and stop. #### Step 2: Domain-Specific Categorization
Once the domain is determined, apply the corresponding sub-rules below. If the domain is "other", skip this step. Explain your reasoning before assigning a label.
- Domain: "writing" - Label 2: Writing Topic - Specify the writing topic in lowercase (e.g., "creative writing", "speech", "LLM prompt", "game", "academic", "social media", "shopping", "formal", "casual", etc.). - If unclear or not explicitly mentioned, use "null". - If none of the predefined topics fit, provide a custom description. - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Creative Writing": Crafting original content, poems, or creative pieces. - "Polishing": Improving clarity, style, and flow. - "Grammar Check": Identifying and correcting grammar mistakes. - "Content Refining": Enhancing word choice, tone, or style. - "Summarization": Condensing text while retaining the core message. - "Expanding": Adding details or elaboration. - "Paraphrasing": Restating content differently. - "Writing Feedback": Providing constructive feedback. - "Translating": Converting text between languages. - "Tone Adjustment": Modifying tone for audience/context. - "Text Summarization": Analyzing meaning, themes, or key points. - If none fit, provide a custom description.
- Domain: "roleplay" - Label 2: Genre - Specify the genre in lowercase (e.g., "sci-fi", "fantasy", "historical", "romance", "mystery", "horror", "action/adventure", "supernatural", etc.). - If multiple genres apply, list them all. - If none fit, provide a custom description. - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Dialogue Generation": Creating natural character conversations. - "Content Creation": Writing from a character’s perspective. - "Question Answering": Responding within character constraints. - "Emotional Simulation": Modeling character-appropriate emotions. - "Decision Support": Providing character-based guidance. - "Skill Training": Simulating expert/practice partners. - "Behavior Simulation": Replicating character behaviors. - "Evaluation & Feedback": Assessing from a character’s perspective. - If none fit, provide a custom description. - Domain: "coding" - Label 2: Programming Language - Specify the language in lowercase (e.g., "python", "java", "c++", "sql", etc.). - If unclear or not mentioned, use "null". - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Code Generation": Creating code from a prompt. - "Debugging": Resolving code errors. - "Code Explanation": Explaining code functionality. - "Code Editing": Modifying existing code. - "Conceptual QA": Answering programming theory questions. - "Code Translation": Converting code between languages. - "Bug Fixing": Fixing specific code issues. - "Testing": Writing/analyzing tests. - "Data Structures & Algorithms": Solving related problems. - "System Design": Designing software architecture. - "Optimization": Improving code efficiency. - "Security": Addressing vulnerabilities. - "Documentation": Writing/improving documentation. - If none fit, provide a custom description. - Domain: "mathematics" - Label 2: Subfield - Specify the subfield in lowercase (e.g., "algebra", "calculus", "geometry", "statistics & probability", etc.). - If unclear, use "null". - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Computation": Performing calculations. - "Proof Writing": Writing formal proofs. - "Conceptual Understanding": Explaining concepts. - "Problem-Solving": Applying techniques to challenges. - "Real-World Application": Connecting math to scenarios. - "Graphing & Visualization": Interpreting/drawing graphs. - "Logical Deduction": Deriving conclusions logically. - If none fit, provide a custom description. - Domain: "reasoning" - Label 2: Reasoning Type - Specify the type in lowercase (e.g., "logical reasoning", "critical thinking", "problem-solving", "analytical thinking", etc.). - If unclear, use "null". - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Logical Deduction": Deriving conclusions logically. - "Puzzle Solving": Solving riddles or brainteasers. - "Argument Evaluation": Analyzing argument validity. - "Scenario Analysis": Considering scenario outcomes. - "Theoretical Reasoning": Exploring abstract concepts. - "Creative Problem-Solving": Finding novel solutions. - "Ethical Reasoning": Evaluating based on ethics. - If none fit, provide a custom description. - Domain: "knowledge" - Label 2: Subject - Specify the subject in lowercase (e.g., "history", "philosophy", "geography", "medicine", "economics", etc.). - If unclear, use "null". - Label 3: Task Type - Choose applicable task types (multiple may apply): - "Fact Recall": Retrieving specific facts. - "Conceptual Understanding": Explaining/interpreting ideas. - "Comparative Analysis": Comparing concepts. - "Causal Reasoning": Examining cause-and-effect. - "Logical Deduction": Applying logic to conclude. - "Hypothetical Thinking": Exploring "what if" scenarios. - "Problem-Solving": Solving practical/theoretical problems. - "Opinion-Based": Providing subjective reasoning. - If none fit, provide a custom description. #### Step 3: Final Output
Return a Python dictionary based on the identified domain and labels:
- For "other": {"domain": "other", "domain_name": "<custom_domain_name>"}
- For specific domains:
{"domain": "<domain>", "task_type": ["<task1>", "<task2>"], "content_type": ["<content1>", "<content2>"]}
 - Replace <domain> with the identified domain (e.g., "writing").
 - Replace "task" with the list of identified task types (or empty list [] if none apply).
 - Replace <content> with the corresponding value (e.g., "poetry", "sci-fi", "history", etc.).
Figure 14: Prompt for initial annotation of tags.
C.1.3 Algorithm for TaxBuilder
The complete TaxBuilder algorithm is presented in Alg. 1.
Algorithm 1 TaxBuilder Algorithm
1:Input: Initial tree $T_{init}$, new nodes $TG_{init}$, LLM-as-Decision-Maker of node insertion DMins
2:Output: Final taxonomy tree $T_{final}$
3:
4:function InsertNode($T,tg$)
[…]
6:     if $d=$<E>" then
7:         return $T$
8:     else if $d=$<S>" then
9:         $T^{c}\leftarrow T^{c}\cup\{tg\}$
10:         return $T$
11:     $\triangleright$ Child level: LLM returns parent node name
12:         $T^{c_{i}}\leftarrow d$
13:         return InsertNode($T^{c_{i}},tg$)
14:     end if
15:end function
16:
17:function TaxBuilder($T_{init}$, $TG_{init}$)
18:     $T_{cur}\leftarrow T_{init}$
19:     for all $tg\in TG_{init}$ do
20:         $T_{cur}\leftarrow$ InsertNode($T_{cur},tg$)
21:     end for
22:     $T_{cur}\leftarrow$ RefinePruneNode($T_{cur}$) $\triangleright$ Node refinement and pruning
23:     $T_{final}\leftarrow$ PruneLayer($T_{cur}$) $\triangleright$ Layer pruing
24:     return $T_{final}$
25:end function
C.1.4 Prompts for TaxBuilder
The prompt for node insertion in the LLM-as-Decision-Maker framework is shown in Fig. 15. The prompt for node refinement and pruning in the LLM-as-Decision-Maker framework is shown in Fig. 16.
Prompt of LLM-as-Decision-Maker (Node Insertion) Your task is to add a new node to a tree-structured classification system. I will give you some [Current Level Nodes] and a [New Node]. You need to determine the next action and respond strictly in the specified format. Before giving the final answer, you must perform brief analysis.
# Task 1
Analyze whether the [New Node] shares identical meaning with any [Current Level Nodes] (including exact matches, case variations, abbreviations, or conceptually equivalent expressions). If identical, return <decision>EXIST</decision> and skip subsequent tasks.
Analyze whether the [New Node] represents a specific category or a classification criterion. If it’s a classification criterion, return <decision>EXIST</decision> and skip subsequent tasks.
# Task 2
Analyze whether the [New Node] doesn’t belong to any subordinate node of [Current Level Nodes] but should be at the same level. If yes, return <decision>ADD</decision> and skip subsequent tasks.
# Task 3
Identify which [Current Level Node] the [New Node] should belong to as a subordinate node. Return the parent node’s name wrapped in <decision></decision>.
## Example
Current Level Nodes: ["Fruits", "Vegetables", "Electronics"]
New Node: Phone
Output: <decision>Electronics</decision>
## New Node
{new_node}
## Current Level Nodes
{current_keys}
## Reference Classification System (for reference only)
{init_tree}
Figure 15: Prompt of LLM-as-Decision-Maker (node insertion).
Prompt of LLM-as-Decision-Maker (Node Refinement and Pruning) Your task is to optimize a given tree structure classification system. I will provide all the next-level nodes under a certain category (referred to as "current layer nodes") and the leaf nodes among them (referred to as "leaf nodes in the current layer"). You need to determine the next steps and return results in the specified format. Before giving the final answer, please provide a brief analysis. Note: Do not modify the names of the original categories, and do not omit any categories from the "current layer nodes" when returning results. Each task analysis should analyze all categories without omission.
If you see that the "current layer nodes" are all classification criteria rather than specific categories, this layer absolutely must not be modified.
### Task 1: Split
Examine all nodes in the "leaf nodes in the current layer" one by one, determine if there are any nodes that are not "single concepts" and split them. If the newly split category already exists, do not return that new category. A "single concept" refers to an independent concept, as opposed to a "combined concept" which is formed by combining multiple concepts. For example, "mythology novel" is a combined concept, consisting of "mythology" and "novel". Return format:
[(old_key1, [new_key1, new_key2]), (old_key2, [new_key1, new_key2])]
### Task 2: Parent-Child Relationship
Analyze all nodes in the "current layer nodes" one by one, determine if there exist parent-child relationships, or if multiple nodes can be merged into a parent node, and return the results. If there are no qualifying situations, do not return anything. Return format:
[(parent_key1, [old_child_key1, old_child_key2]), (parent_key2, [old_child_key1, old_child_key2])]
### Task 3: Delete
Examine all nodes in the "current layer nodes" one by one, identify meaningless categories (such as null values, blanks, etc.) or extremely rare categories in the real world, and delete them. Return format:
[old_key_to_delete1, old_key_to_delete2]
### Task 4: Merge
Analyze all nodes in the "current layer nodes" one by one, identify categories with consistent concepts and merge them. Return format:
[(new_key1, [old_key_to_merge1, old_key_to_merge2]), (new_key2, [old_key_to_merge1, old_key_to_merge2])]
### Task 5: Keep
Preserve the remaining nodes. Return format:
[old_key1, old_key2]
## Output Format
Provide a brief analysis for each task and give the results after the task is executed.
After completing all tasks, format the final result as a Python dictionary wrapped in <decision></decision> tags, in the following format, where all keys with old_ must have appeared in the "current layer nodes":
<decision> {
"split": [(old_key1, [new_key1, new_key2]), (old_key2, [new_key1, new_key2])],
"reparent": [(parent_key1, [old_child_key1, old_child_key2]), (parent_key2, [old_child_key1, old_child_key2])],
"delete": [old_key_to_delete1, old_key_to_delete2],
"merge": [(new_key1, [old_key_to_merge1, old_key_to_merge2]), (new_key2, [old_key_to_merge1, old_key_to_merge2])],
"keep": [old_key1, old_key2]
} </decision>
If there is no operation for a certain task, do not display that item in the final output.
## Current Layer Nodes
{current_keys}
## Leaf Nodes in the Current Layer
{leaf_keys}
Figure 16: Prompt of LLM-as-Decision-Maker (node refinement and pruning).
C.2 RealMix
C.2.1 Seed Data
The prompt for labeling the domain is shown in Fig.17.
The prompts for query tagging are shown in Fig.18, Fig.19, Fig.20, Fig.21, Fig.22, and Fig.23.
The prompts for quality labeling are shown in Fig.24, Fig.25, Fig.26, Fig.27, Fig.28, and Fig. 29.
The seed data is available in our dataset repository.
Prompt of Domain Annotation You are tasked with categorizing a given query into one of the following six categories:
1. **roleplay**
2. **coding**
3. **mathematics**
4. **reasoning**
5. **knowledge**
6. **writing**
7. **other**
For each query provided, determine the most appropriate category and output the result in lowcase enclosed within <domain> and </domain> tags.
**Example**:
Query: "How do I write a compelling essay introduction?"
Output: <domain>roleplay</domain>
**Now, analyze the following query**:
<|begin_of_query|>
{query}
<|end_of_query|>
Figure 17: Prompt of domain annotation.
Prompt of Query Tags Annotation (writing) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system.
### Classification System:
{taxonomy}
### Rules for Tagging:
1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags.
2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes.
3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category.
4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags.
### Example:
Query: "I need help brainstorming ideas for a fantasy story involving a hidden kingdom and magical creatures."
Output: <tags>{{"Creative Stages": ["Conceptualization Stage", "Brainstorming", "World Building", "Creative Exploration", "Content Conceptualization"], "Writing Domains": ["Literary Writing", "Fiction", "Fantasy", "Story Creation"], "Styles": ["Other"]}}</tags>
### Now, analyze the following query:
<|begin_of_query|>
{query}
<|end_of_query|>
Figure 18: Prompt of query tags annotation (writing).
Prompt of Query Tags Annotation (roleplay) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system.
### Classification System:
{taxonomy}
### Rules for Tagging:
1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags.
2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes.
3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category.
4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags.
### Example:
Query: "Write a script for a medieval fantasy story involving a knight’s adventure and a magical encounter, with a humorous twist."
Output: <tags>{{"Theme Types": ["Fantasy", "Medieval", "Magic", "Comedy", "Humorous Comedy"], "Task Types": ["Creative", "Scriptwriting"], "Style Types": ["Humorous Style", "Light Humor", "Narrative"]}}</tags>
### Now, analyze the following query:
<|begin_of_query|>
{query}
<|end_of_query|>
Figure 19: Prompt of query tags annotation (roleplay).
Prompt of Query Tags Annotation (knowledge) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system.
### Classification System:
{taxonomy}
### Rules for Tagging:
1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags.
2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes.
3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category.
4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags.
### Example:
Query: "How can I analyze data from a physics experiment on thermodynamics and present it effectively?"
Output: <tags>{{"Disciplinary Fields": ["Natural Sciences", "Physics", "Thermodynamics"], "Cognitive Levels": ["Applied Analysis"], "Task Types": ["Information Processing", "Data Analysis", "Content Production", "Content Generation"]}}</tags>
### Now, analyze the following query:
<|begin_of_query|>
{query}
<|end_of_query|>
Figure 20: Prompt of query tags annotation (knowledge).
Prompt of Query Tags Annotation (coding) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system.
### Classification System:
{taxonomy}
### Rules for Tagging:
1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags.
2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes.
3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category.
4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags.
### Example:
Query: "How can I optimize a Python script that processes large datasets and visualizes the results?"
Output: <tags>{{"Task Types": ["Data Processing", "Data Analysis", "Data Visualization", "Debugging and Optimization", "Optimization"], "Technical Domains": ["Data Science", "Data Analysis Methods", "Data Visualization Tools"], "Programming Languages": ["General-purpose Languages", "Python"]}}</tags>
### Now, analyze the following query:
<|begin_of_query|>
{query}
<|end_of_query|>
Figure 21: Prompt of query tags annotation (coding).
Prompt of Query Tags Annotation (mathematics) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system.
### Classification System:
{taxonomy}
### Rules for Tagging:
1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags.
2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes.
3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category.
4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags.
### Example:
Query: "How do I use calculus to model the growth of a population over time and graph the results?"
Output: <tags>{{"Mathematical Subfields": ["Analysis", "Calculus", "Applied Mathematics"], "Task Types": ["Problem Solving", "Modeling", "Visualization", "Function Graphing"]}}</tags>
### Now, analyze the following query:
<|begin_of_query|>
{query}
<|end_of_query|>
Figure 22: Prompt of query tags annotation (mathematics).
Prompt of Query Tags Annotation (reasoning) You are tasked with categorizing a given query into multiple tags based on the following hierarchical classification system.
### Classification System:
{taxonomy}
### Rules for Tagging:
1. **Hierarchy Rule**: If a query matches both a parent node and its child nodes, include both in the tags.
2. **Multiple Matches**: If a query matches multiple nodes at the same level, include all matching nodes.
3. **No Match**: If a query does not match any nodes under the second-level main categories, assign the tag "Other" to that category.
4. **Output Format**: The final output must be enclosed within ‘<tags>‘ and ‘</tags>‘ tags, and the tags should be provided as a JSON object where the keys are the basis for classification and the values are lists of matching tags.
### Example:
Query: "How can I determine the cause of an event based on multiple contributing factors?"
Output: <tags>{{"Reasoning Methods": ["Multi-factor Attribution"], "Application Domains": ["Other"], "Thinking Modes": ["Analytical"], "Task Types": ["Reasoning"]}}</tags>
### Now, analyze the following query:
<|begin_of_query|>
{query}
<|end_of_query|>
Figure 23: Prompt of query tags annotation (reasoning).
Prompt of Quality Annotation (writing) ## Task
Your task is to assess the quality of a given query based on the checklists outlined below. For each checklist, provide a clear explanation of your reasoning before assigning "Yes" or "No". If your answer is "Borderline", please answer "Yes" for this checklist.
## Checklists
Determine whether the provided question meets the following criteria. Return a Python array listing the numbers of all satisfied criteria:
1. **Clarity**: Is the question clear and well-defined?
2. **Completeness**: Does the question provide enough information for the LLM to answer the question?
3. **Complexity**: Does the question have enough depth and challenge beyond simple fact recall?
4. **Real-World Application**: Is this writing task something that would be proposed in the real world?
5. **Professionalism**: Does it require professional capabilities or professional knowledge?
6. **Originality:** Does the question encourage or require originality?
7. **User’s Requirements**: Does the user have clear, detailed, or unique requests that need to be considered in the response?
For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘.
## Final Output
For each question provided, return a Python dictionary in the following format:
“‘python
Final Labels: {{"question_quality": [1, 3, 4]}}
## The Query You Should Process
<|begin_of_query|>
{query}
<|end_of_query|>
Figure 24: Prompt of quality annotation (writing).
Prompt of Quality Annotation (roleplay) ## Task
Your task is to assess the quality of a given query based on the checklists outlined below. For each checklist, provide a clear explanation of your reasoning before assigning "Yes" or "No". If your answer is "Borderline", please answer "Yes" for this checklist.
## Checklists
Determine whether the provided question meets the following criteria. Return a Python array listing the numbers of all satisfied criteria:
1. **Clarity**: Is the question clear and well-defined?
2. **Completeness**: Does the question provide enough information for the LLM to answer the question?
3. **Complexity**: Does it involve in-depth understanding of any role-playing content, such as the psychology, characterization, and world-building of characters?
4. **Real-World Application**: Is this role-playing task something that would be proposed in the real world?
5. **Interactivity**: Does the question encourage meaningful interactions between characters, rather than single character?
6. **Engagement**: Does the task motivate active participation and emotional involvement from the audience or participants?
7. **Creativity:** Does it have creativity and novelty, or does solving it require creativity?
For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘.
## Final Output
For each question provided, return a Python dictionary in the following format:
“‘python
Final Labels: {{"question_quality": [1, 3, 4]}}
## The Query You Should Process
<|begin_of_query|>
{query}
<|end_of_query|>
Figure 25: Prompt of quality annotation (roleplay).
Prompt of Quality Annotation (knowledge) ## Task
Your task is to assess the quality of a given query based on the checklists outlined below. For each checklist, provide a clear explanation of your reasoning before assigning "Yes" or "No". If your answer is "Borderline", please answer "Yes" for this checklist.
## Checklists
Determine whether the provided question meets the following criteria. Return a Python array listing the numbers of all satisfied criteria:
1. **Clarity**: Is the question clear and well-defined?
2. **Completeness**: Does the question provide enough information for the LLM to answer the question?
3. **Complexity**: Does the question have enough depth and challenge beyond simple fact recall?
4. **Real-World Application**: Is the question something that would be encountered in real-world?
5. **Depth of Knowledge**: Does the question require deep expertise in the subject instead of just memory?
6. **Cross-Disciplinary**: Does the question involve cross-disciplinary aspects?
7. **Open-Endedness.**: Does the question encourage open-ended responses rather than simple “yes” or “no” answers, promoting deeper thinking?
For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘.
## Final Output
For each question provided, return a Python dictionary in the following format:
“‘python
Final Labels: {{"question_quality": [1, 3, 4]}}
## The Query You Should Process
<|begin_of_query|>
{query}
<|end_of_query|>
Figure 26: Prompt of quality annotation (knowledge).
Prompt of Quality Annotation (coding) ## Task
Your task is to assess the quality of a given query based on the checklists outlined below. For each checklist, provide a clear explanation of your reasoning before assigning "Yes" or "No". If your answer is "Borderline", please answer "No" for this checklist.
## Checklists
Determine whether the provided question meets the following criteria. Return a Python array listing the numbers of all satisfied criteria:
1. **Clarity**: Is the question clear and well-defined?
2. **Completeness**: Does the question provide enough information for the LLM to answer the question?
3. **Complexity**: Does it involve multiple components, layers, or nuance?
4. **Real-World Application**: Is the question something that would be encountered in real-world development?
5. **Problem-Solving**: Does it require active problem-solving beyond simple and superficial script or fact recall?
6. **Domain-Specific Expertise**: Does the question require in-depth knowledge of at least one specific area of programming?
7. **Specified Requirements**: Does it specify particular requirements, such as execution time, space constraints, specific programming language, tools, packages, etc.?
For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘.
## Final Output
For each question provided, return a Python dictionary in the following format:
“‘python
Final Labels: {{"question_quality": [1, 3, 4]}}
## The Query You Should Process
<|begin_of_query|>
{query}
<|end_of_query|>
Figure 27: Prompt of quality annotation (coding).
Prompt of Quality Annotation (mathematics) ## Task
Your task is to assess the query of a given question based on the rules outlined below. For each rule, provide a clear explanation of your reasoning before assigning a label.
## Rule
Determine whether the provided query meets the following criteria. Return a Python array listing the numbers of all satisfied criteria:
1. **Clarity**: Is the question clear and well-defined?
2. **Completeness**: Does the question provide enough information for the LLM to answer the question?
3. **Complexity**: Does it involve multiple steps, analysis, or reasoning instead of simple concept memorization and numerical calculation?
4. **Real-World Application**: Is the question something that would be encountered in real-world?
5. **Problem-Solving**: Does it test the ability to apply math in some scenarios?
6. **Rigorous Logic**: Does it involve content such as theorem derivation and formula understanding, which require rigorous logical abilities?
7. **Creativity:** Does it have creativity and novelty, or does solving it require creativity?
For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘.
## Final Output
For each question provided, return a Python dictionary in the following format:
“‘python
Final Labels: {{"question_quality": [1, 3, 4]}}
## The Query You Should Process
<|begin_of_query|>
{query}
<|end_of_query|>
Figure 28: Prompt of quality annotation (mathematics).
Prompt of Quality Annotation (reasoning) ## Task
Your task is to assess the query of a given question based on the rules outlined below. For each rule, provide a clear explanation of your reasoning before assigning a label.
## Rule
Determine whether the provided query meets the following criteria. Return a Python array listing the numbers of all satisfied criteria:
1. **Clarity**: Is the question clear and well-defined?
2. **Completeness**: Does the question provide enough information for the LLM to answer the question?
3. **Complexity**: Does it involve multiple steps, analysis, or reasoning instead of simple concept memorization?
4. **Real-World Application**: Is the question something that would be encountered in real-world?
5. **Problem-Solving**: Does it require devising a solution or strategy?
6. **Deep Thinking**: Does it require deep reasoning and consideration of multiple factors?
7. **Novelty:** Does the question present a unique or unusual scenario that the LLM is unlikely to have encountered before?
For example, if the question meets Clarity, Completeness, and Real-World Application, return ‘[1, 2, 4]‘.
## Final Output
For each question provided, return a Python dictionary in the following format:
“‘python
Final Labels: {{"question_quality": [1, 3, 4]}}
## The Query You Should Process
<|begin_of_query|>
{query}
<|end_of_query|>
Figure 29: Prompt of quality annotation (reasoning).
C.2.2 Synthesize New Queries
The prompt of RealMix is shown in Fig. 30
Prompt of RealMix I will provide three real user questions and a reference question’s type tags. Your task is to create a new question in the {domain} domain with the same tags as reference question. To ensure the question feels authentic, you should utilize real-world details drawn from the three real user questions.
### You MUST follow these steps to generate the new question:
#### Task1: Real-world Details Selection:
Select a few real-world details (such as the real people, objects, scenes, settings, and any other details mentioned) from three real user questions that fit the given tags. If there are no suitable details, return [[I cannot generate a question based on the provided real user questions.]] and stop.
#### Task2: New Question Generation:
1. Although you should use details from real user questions, you must not mention the real user question in the new question.
2. The new question should be complex and challenging, requiring deep understanding and analysis of the subject. The length of the question should be at least as long as the reference question but should not be overly simplistic or repetitive. The question should be singular, not a multi-task question.
3. The new question must be **completely self-contained**, so that others can answer it without any additional information.
4. Analyze how to create the new question with chosen real-world details and provided tags. While multiple tags are available, the newly generated question only needs to align with some of them, not all. Even if the original question already fits, generate a different version.
### Output Format:
[Anylysis]: You should first complete the anylysis of task1 and task2 here.
[Question]: Summarize the newly generated question in the following format: <new_query>Insert the final new question here.</new_query>
<|begin_of_reference_query|>
{reference_query}
<|end_of_reference_query|>
<|begin_of_reference_query_type_tags|>
{type_tags}
<|end_of_reference_query_type_tags|>
<|begin_of_real_user_query_1|>
{query1}
<|end_of_real_user_query_1|>
<|begin_of_real_user_query_2|>
{query2}
<|end_of_real_user_query_2|>
<|begin_of_real_user_query_3|>
{query3}
<|end_of_real_user_query_3|>
Figure 30: Prompt of RealMix.
C.3 Detailed Description of Visualization and Analysis Toolkits
Visualization. Two visualization tools are available: one organizes results by model, enabling side-by-side comparisons across multiple models; the other organizes results by query type, allowing users to examine model rankings under specific query tags.
Analysis. We have implemented four analysis tools, with more under development.
  • Automatic Weakness Reporter presents a list of notable strengths and weaknesses detected in each model.
  • Automatic Weakness Analyzer generates natural language summaries of a model’s strengths and weaknesses.
  • Failure Mode Explorer identifies whether a model’s failure in a subtask is systematic (low ranking variance among sibling nodes) or sporadic (high variance).
  • Automatic Difference Reporter compares multiple models and highlights significant anomalies between their performances in a structured list.
C.4 Feedbacker-D-V0
C.4.1 Query Quality: Human Evaluation
We conduct a comprehensive questionnaire study to evaluate the quality and human-likeness of automatically generated questions compared to original real-user questions. Five domain experts (four Master’s holders and one PhD holder) participate in the questionnaire. The assessment protocol, shown in Fig. 31, present participants with 20 question pairs (one generated and one original) in randomized order to prevent positional bias.
Query Quality and Reality Evaluation Questionnaire Dear Participant, We are conducting a study on question generation for large language model. The attached Excel file contains 20 rows, each presenting two questions generated by different strategies. Please: 1. Assess the quality of both questions (poor quality indicators include: oversimplification, incompleteness, or unclear phrasing). 2. Evaluate whether each question appears to be human-authored. Rating Scheme: For quality/authenticity: 1 = Question 1 superior 2 = Question 2 superior 3 = Both inadequate 4 = Both excellent
Figure 31: Questionnaire protocol for comparative evaluation.
Tab. 6 presents the detailed evaluation outcomes across all raters. Our generated questions demonstrate superior quality in 49% of cases compared to just 9% where original questions are preferred, with 42% of cases rated as ties. For authenticity assessment, generated questions are preferred in 27% of cases versus 21% for originals, with a majority (52%) considered equally authentic. Notably, 37 out of 100 cases are rated as having both excellent quality, while 47 cases are judged equally authentic.
Table 6: Query quality and reality evaluation results by human rater.
Rating Rater 1 Rater 2 Rater 3 Rater 4 Rater 5 Total
Quality/Reality Quality/Reality Quality/Reality Quality/Reality Quality/Reality Quality/Reality
1 (Generated better) 11/5 15/8 10/3 7/7 6/4 49/27
2 (Original better) 0/4 4/10 4/7 0/0 1/0 9/21
3 (Both inadequate) 1/1 0/0 2/3 2/1 0/0 5/5
4 (Both excellent) 8/10 1/2 4/7 11/12 13/16 37/47
C.4.2 Query Quality: Data Contamination
Following Auto-Arena [19], we employ two methods to detect test data contamination. The first is the string match metric, as used in GPT-4 [71], which flags a data point as contaminated if any of three randomly sampled 50-character substrings (or the entire string, if shorter) from the evaluation sample appears as a substring in the training set. The second method, inspired by Platypus [72], uses sentence embedding similarity: a question is considered contaminated if its BERT-Large [73] embedding has a cosine similarity above 0.8 with any training item, enabling detection of paraphrased overlaps.
Applying these methods, we find a contamination rate of 1.41% using the substring match metric, and a substantially lower rate of 0.21% using the embedding similarity metric. These results indicate that our approach is largely robust to data leakage and that potential contamination is minimal.
C.4.3 Examples of Feedbacker-D-V0
We present several samples of Feedbacker-D-V0 in Tab. 7.
Table 7: Examples from different domains in Feedbacker-D-V0.
Domain Sample
Writing Translate and adapt the following mission dialogue into French, ensuring it resonates with a native French speaker. The dialogue should maintain its dramatic tone while incorporating cultural nuances and idiomatic expressions suitable for a fantasy RPG setting. Additionally, optimize the sentence structure to enhance flow and engagement. "Listen carefully, adventurer. The ancient temple lies hidden beyond the Whispering Woods. Within its walls, you will find the Earth Crystal, a source of immense power. Use it wisely, for it can either restore balance or bring about destruction. Be wary of the Guardians; they will test your worth. Prove yourself, and the crystal shall be yours."
Roleplay Please pretend to be my deceased grandfather, who was a gentle and wise artist from the small coastal town of Lestupsk. He used to live in a house near the old fortress on the hill and would often take me to ride the drowsy trams that clanked through our town. He taught me how to draw and always shared stories about his friend, Vanya, another local artist. I miss him so much, especially when I am trying to fall asleep. We begin now. Hello grandpa, I miss you so much! I am feeling very tired and am having trouble falling asleep. Can you please tell me one of those special stories about Vanya and the old trams, and maybe show me how to sketch the view of the fortress from our window? Your stories always made me feel better and helped me sleep.
Knowledge Analyze how the infrared absorption spectrum of a greenhouse gas (e.g., CO2 or methane) impacts the heat transfer efficiency in a thermodynamic system such as a gas turbine or a heat exchanger. Propose strategies to counteract any detrimental effects on system performance caused by such radiative properties while maintaining environmental compliance.
Coding In Java, which method is commonly used to remove quotes from the beginning and end of a string? a) ‘trim()‘ b) ‘replaceAll(""̈, "")‘ c) ‘substring(1, str.length() - 1)‘ d) ‘strip()‘ e) None of the above
Mathematics Let $F=A^{\prime}B^{\prime}C^{\prime}+A^{\prime}B^{\prime}C+A^{\prime}BC^{\prime}+AB^{\prime}C^{\prime}+ABC^{\prime}$. Using Boolean algebraic laws, simplify the expression $F$ to its minimal form. Identify which laws or manipulations you use at each step to arrive at the solution.
Reasoning A teacher places a book inside a drawer of her desk in the classroom. She then locks the drawer and places the key inside a box on her shelf. After school, the teacher moves the box to the staff room and leaves it on a table. Later, a janitor takes the box from the staff room and places it inside a cabinet in the storage room. Where is the key, and where is the book?
C.4.4 Feedbacker-HPD
To validate our evaluation methods, we require a robust human preference dataset. We construct this dataset from seed data, which is well-suited for our purposes as it contains two responses and a corresponding human preference label for each query. We begin by randomly sampling 650 queries rated 6 or 7 from the seed data, followed by manual quality control, yielding a final set of 636 high-quality queries.
Although the seed data includes human preference labels, prior work [15] has raised concerns about their reliability due to variability in human judgment. To enhance label accuracy, we enlisted two additional graduate-level annotators and determined the final preference label via majority voting, thereby mitigating individual annotator bias. Our annotation interface is illustrated in Fig. 32.
Since our questions are all compliant and commonly seen in real-world scenarios, they do not pose significant potential participant risks. Additionally, we pay 300 yuan per day for each participant, and we ultimately spent 1,200 yuan.
The distribution of Feedbacker-HPD is shown in Fig. 33. Similar to Feedbacker-D-V0, these queries span six domains: writing, roleplay, knowledge, coding, mathematics, and reasoning.
Figure 32: The annotation interface used for labeling the human preference dataset.
Figure 33: Domain distribution of Feedbacker-HPD.
Appendix D Evaluation Methods
Example (1) of Pairwise Evaluation (prompt and thinking process) …
[[A]]
**Explanation:**
- **Correctness:** Assistant A’s answer directly addresses the user’s requirement to track the *number of open tabs* on the same domain using ‘localStorage‘. It provides a straightforward ‘TabTracker‘ class with methods to increment/decrement the count and store it in ‘localStorage‘. While it lacks automatic triggering of decrement on tab closure (relying on manual calls in the example), it correctly structures the core logic for counting tabs. The problem does not explicitly require automatic event handling, so the code is technically correct for the stated goal.
- **Helpfulness:** The code is concise and focused on the core requirement. The explanation clearly describes how each method works and how to use the class. The example usage demonstrates the basic workflow, making it easy to integrate into a project.
- **Assistant B’s Issues:** While B’s approach is more detailed (tracking individual tabs with IDs and URLs), it diverges from the user’s request to track *open tabs as a count*. B’s implementation requires managing individual tabs and resolving complex edge cases (e.g., identifying which tab closed), which complicates the task. Additionally, B’s ‘_handleTabClose‘ method admits it cannot reliably remove tabs, making the solution incomplete for the stated problem.
- **Key Flaw in B’s Answer:** The user asked for a *count* of open tabs, not detailed tracking of individual tabs. B’s answer overcomplicates the solution by focusing on tab metadata rather than the numeric count.
**Conclusion:** Assistant A’s answer is better because it aligns directly with the user’s requirement, provides a functional core, and is simpler and more relevant. Assistant B’s answer, while thorough, solves a different problem (tracking individual tabs) and introduces unnecessary complexity.
Figure 34: Example (1) of Pairwise Evaluation.
Example (2) of Pairwise Evaluation (prompt and thinking process) …
[[B]]
**Explanation:**
Both assistants provided valid HTML documents with step-by-step instructions for jump-starting a car. However, **Assistant B’s answer is better** for the following reasons:
1. **Self-Contained Diagram:**
- **Assistant B** uses an embedded SVG diagram directly in the HTML, fulfilling the user’s requirement for an "all in one document." The SVG illustrates the connection process without relying on an external image file.
- **Assistant A** references an external image (‘car_battery_jumpstart.png‘), which requires the user to provide or replace the image path. This makes the document incomplete as-is.
2. **Step-by-Step Clarity:**
- **Assistant B** breaks down the steps into a more granular 8-step process, explicitly separating each cable connection into distinct steps (e.g., connecting each end of the red and black cables separately). This enhances clarity for users unfamiliar with the process.
- **Assistant A** combines some steps (e.g., all cable connections under step 3 with sub-steps), which is concise but slightly less explicit.
3. **Safety and Correctness:**
- Both answers include correct procedural steps (e.g., grounding the negative cable on the dead car’s metal surface instead of the terminal).
- **Assistant B** adds a final step (step 8) explicitly detailing the order to remove cables, which is important for safety.
4. **HTML Structure:**
- **Assistant B’s** use of SVG ensures the diagram is part of the HTML document, whereas **Assistant A’s** reliance on an external image breaks the "all in one" requirement.
**Verdict:** While both answers are correct, **Assistant B’s** self-contained implementation of the diagram and clearer step breakdown make it the superior choice.
Figure 35: Example (2) of Pairwise Evaluation.